# Exact overlaps for ”all” integrable matrix product states of rational spin chains

### v1

@@ ACKNOWLEDGMENTS
$\mathbf{F}^{(n)}(u)=\left[\tilde{\mathbf{K}}_{2,-1}^{(k)}(u)\right]^{-1}\left[\tilde{\mathbf{K}}_{-1,2}^{(k)}(u)-\tilde{\mathbf{K}}_{-1,-1}^{(k)}(u)\left[\tilde{\mathbf{K}}_{2,-1}^{(k)}(u)\right]^{-1}\tilde{\mathbf{K}}_{2,2}^{(k)}(u)\right].$ (S-54)
S-IV.1.2 $Y(\mathfrak{o}_{4k+2},\mathfrak{gl}_{2n+1})$ (achiral)
We define the nested $K$-matrices $\tilde{\mathbf{K}}^{(s)}$ for $s=1,\dots,k$ with the recursion (S-50) and initial condition $\tilde{\mathbf{K}}^{(1)}\equiv\mathbf{K}$. We also define the $Y^{-}(2)$ $K$-matrices $\mathbf{k}^{(s)}$ the deformed ones $\tilde{\mathbf{k}}^{(s)}$ for $s=1,\dots,k$ with (S-52). The G-operators $\mathbf{G}^{(1)},\dots,\mathbf{G}^{(n-1)}$ are defined with (S-37) for $s=1,\dots,k$. The remaining $G$-operator is
$\mathbf{G}^{(n)}(u)=\tilde{\mathbf{K}}_{1,1}^{(k)}(u)-\sum_{\alpha,\beta=2}^{3}\tilde{\mathbf{K}}_{1,-\alpha}^{(k)}(u)\widehat{\mathbf{K}}_{-\alpha,\beta}^{(k)}(u)\tilde{\mathbf{K}}_{\beta,1}^{(k)}(u).$ (S-55)
@@
The G-operators are the same as for $Y(\mathfrak{o}_{2n+1},\mathfrak{o}_{n}\oplus\mathfrak{o}_{n+1})$ and the difference is only in the last step. We have a series of $K$-matrices $\mathbf{K}^{(s)}$ for $s=1,\dots,n-1$ with the recursion (S-40). We also define the G-operators $\mathbf{G}^{(s)}$ for $s=1,\dots,n-1$ with (S-41). We also need to define the remaining $G$-operators $\mathbf{G}^{(n)}(u),\mathbf{G}^{(n+1)}(u)$ as
$\begin{split}\mathbf{G}^{(n)}(u)&=\mathbf{K}_{1,-1}^{(n-1)}(u)-\mathbf{K}_{1,-2}^{(n-1)}(u)\left[\mathbf{K}_{2,-2}^{(n-1)}(u)\right]^{-1}\mathbf{K}_{2,-1}^{(n-1)}(u),\\ \mathbf{G}^{(n+1)}(u)&=\mathbf{K}_{-1,1}^{(n-1)}(u)-\mathbf{K}_{-1,-2}^{(n-1)}(u)\left[\mathbf{K}_{2,-2}^{(n-1)}(u)\right]^{-1}\mathbf{K}_{2,1}^{(n-1)}(u).\end{split}$ (S-56)
@@
$\mathbf{G}^{(n)}(u)=\mathbf{K}_{1,1}^{(n-1)}(u)-\mathbf{K}_{1,-2}^{(n-1)}(u)\left[\mathbf{K}_{2,-2}^{(n-1)}(u)\right]^{-1}\mathbf{K}_{2,1}^{(n-1)}(u).$ (S-57)
@@
We continue with general $n$ and $M$. The G-operators are the same as for $Y(\mathfrak{o}_{2n+1},\mathfrak{o}_{M}\oplus\mathfrak{o}_{2n-M+1})$ and the difference is only in the last step. We have a series of $K$-matrices $\mathbf{K}^{(s)}$ for $s=1,\dots,M+1$ with the recursion (S-40). We also define the G-operators $\mathbf{G}^{(s)}$ for $s=1,\dots,M$ with (S-41). We also need to define the remaining $G$-operators $\mathbf{G}^{(M+1)}(u),\mathbf{G}^{(M+2)}(u),\dots,\mathbf{G}^{(n+1)}(u)$ for the chiral and $\mathbf{G}^{(M+1)}(u),\mathbf{G}^{(M+2)}(u),\dots,\mathbf{G}^{(n)}(u)$ for achiral cases. When $n-M$ is even the pair structure is chiral and the pair structure is odd when $n-M$ is odd.
@@
$\begin{split}\mathbf{K}_{a,b}(u)&=\sum_{c=-\tilde{n}}^{\tilde{n}}\mathbf{L}_{-c,-a}(u-\kappa_{N-2M})\mathbf{L}_{c,b}(-u)\\ \tilde{\mathbf{K}}_{a,b}^{(1)}(u)&=\mathbf{K}_{a,b}(u)+\\ &+u\sum_{l=1}^{k}\left(\mathbf{L}_{-(n-2l+1),-a}(u-\kappa_{N-2M})\mathbf{L}_{-n+2l-2,b}(-u)-\mathbf{L}_{-(n-2l+2),-a}(u-\kappa_{N-2M})\mathbf{L}_{-n+2l-1,b}(-u)\right).\end{split}$ (S-58)
where $\tilde{n}=2k=n-M$. We continue the series of the nested $K$-matrices $\tilde{\mathbf{K}}_{a,b}^{(s)}(u)$ for $s=2,\dots,k$ as (S-50) and define the $Y^{-}(2)$ $K$-matrices $\mathbf{k}^{(s)}$ and the deformed ones $\tilde{\mathbf{k}}^{(s)}$ for $s=1,\dots,k$ with (S-52) (we replace $n\to\tilde{n}$ , $\mathbf{K}^{(s)}\to\tilde{\mathbf{K}}^{(s)}$ and $\kappa_{N-4s+4}\to\kappa_{N-2M-4s+4}$). The G-operators $\mathbf{G}^{(M+1)},\dots,\mathbf{G}^{(n)}$ with
$\begin{split}\mathbf{G}^{(M+2s-1)}(u)&=\tilde{\mathbf{k}}_{1,1}^{(s)}(u),\\ \mathbf{G}^{(M+2s)}(z)&=\tilde{\mathbf{k}}_{2,2}^{(s)}(u)-\tilde{\mathbf{k}}_{2,1}^{(s)}(u)\left[\tilde{\mathbf{k}}_{1,1}^{(s)}(u)\right]^{-1}\tilde{\mathbf{k}}_{1,2}^{(s)}(u),\end{split}$ (S-59)
for $s=1,\dots,k$. The F-operators are defined as (S-53) and (S-54).
If $n-M$ is odd, i.e. $n-M=2k+1$ ($k\in\mathbb{Z}_{+}$) we have to deform the K-matrix $\mathbf{K}^{(M+1)}$ as (S-58) where $\tilde{n}=2k+1=n-M$. We continue the series of the nested $K$-matrices $\tilde{\mathbf{K}}_{a,b}^{(s)}(u)$ for $s=2,\dots,k$ as (S-50). We define the $Y^{-}(2)$ $K$-matrices $\mathbf{k}^{(s)}$ and the deformed ones $\tilde{\mathbf{k}}^{(s)}$ for $s=1,\dots,k$ with (S-52) (we replace $n\to\tilde{n}$ , $\mathbf{K}^{(s)}\to\tilde{\mathbf{K}}^{(s)}$ and $\kappa_{N-4s+4}\to\kappa_{N-2M-4s+4}$). The G-operators $\mathbf{G}^{(M+1)},\dots,\mathbf{G}^{(n-1)}$ with (S-59) for $s=1,\dots,k$. The remaining $G$-operator $\mathbf{G}^{(n)}$ is defined as (S-55).
@@
In the $\mathfrak{g}=\mathfrak{o}_{2n+1}$ case there exists only chiral pair structure. We introduce $n+1$ G-operators $\mathbf{G}^{(k)}(u)$ for $k=1,\dots,n+1$ and we introduce the $F$-operators as (S-23) and the final $\tilde{\mathcal{F}}$ functions defined as
~$\tilde{\mathcal{F}}_{j}^{(k)}(u)=\begin{cases}\mathcal{F}_{j}^{(k)}(-iu-\frac{1}{4}+\frac{n-k}{2})\sqrt{\frac{u^{2}}{u^{2}+1/4}},&k=1,\dots,n-1,\\ \mathcal{F}_{j}^{(n)}(-iu-\frac{1}{4})\frac{u-i/4}{u-i/2}\sqrt{\frac{u^{2}}{u^{2}+1/16}},&k=n.\end{cases}$ (S-60)
@@
For $Y(\mathfrak{o}_{2n+1},\mathfrak{o}_{n}\oplus\mathfrak{o}_{n+1})$ we have a series of $K$-matrices $\mathbf{K}^{(k)}$ for $k=1,\dots,n$ with the recursion (S-40) with the initial condition $\mathbf{K}^{(1)}(u)\equiv\mathbf{K}(u)$. We also define the G-operators $\mathbf{G}^{(k)}$ for $k=1,\dots,n$ with (S-41) and $\mathbf{G}^{(n+1)}$ with (S-42).
@@
For general $M$ we combine the methods of the previous sections. We have a series of $K$-matrices $\mathbf{K}^{(k)}$ for $k=1,\dots,M+1$ with the recursion (S-40) and initial condition $\mathbf{K}^{(1)}(u)\equiv\mathbf{K}(u)$. We also define the G-operators $\mathbf{G}^{(k)}$ for $k=1,\dots,M$ with (S-41). We also need to define the remaining $G$-operators $\mathbf{G}^{(M+1)}(u),\mathbf{G}^{(M+2)}(u),\dots,\mathbf{G}^{(n+1)}(u)$. Their definitions depend on whether $n-M$ is even or odd.
If $n-M$ is odd, i.e. $n-M=2k+1$ ($k\in\mathbb{N}$) we have to deform the K-matrix $\mathbf{K}^{(M+1)}$ with (S-58) where $\tilde{n}=2k+1=n-M$. We define the nested $K$-matrices $\tilde{\mathbf{K}}_{a,b}^{(s)}(u)$ for $s=2,\dots,k+1$ with (S-50), the $Y^{-}(2)$ $K$-matrices $\mathbf{k}^{(s)}$ and the deformed ones $\tilde{\mathbf{k}}^{(s)}$ for $s=1,\dots,k$ with (S-52) (we replace $n\to\tilde{n}$ , $\mathbf{K}^{(s)}\to\tilde{\mathbf{K}}^{(s)}$ and $\kappa_{N-4s+4}\to\kappa_{N-2M-4s+4}$). We can also define the G-operators $\mathbf{G}^{(r)}$ for $r=M+1,\dots,n-1$ as (S-59) for $s=1,\dots,k$. We also have to deform the K-matrix $\tilde{\mathbf{K}}_{a,b}^{(k+1)}(u)$ as
$\begin{split}\tilde{\mathbf{K}}_{a,b}^{(k+1)}(u)&=\sum_{c=-1}^{1}\mathbf{L}_{-c,-a}(u-\kappa_{3})\mathbf{L}_{c,b}(-u)\\ \check{\mathbf{K}}_{a,b}^{(k+1)}(u)&=\tilde{\mathbf{K}}_{a,b}^{(k+1)}(u)-\frac{1}{2}u(u+\frac{1}{4})\mathbf{L}_{-1,-a}(u-\kappa_{3})\mathbf{L}_{-1,b}(-u).\\ &+u\left(\mathbf{L}_{0,-a}(u-\kappa_{3})\mathbf{L}_{-1,b}(-u)-\mathbf{L}_{-1,-a}(u-\kappa_{3})\mathbf{L}_{0,b}(-u)\right).\end{split}$ (S-61)
and the remaining $G$-operators $\mathbf{G}^{(n)}$, $\mathbf{G}^{(n+1)}$ as
$\begin{split}\mathbf{G}^{(n)}(u)&=\check{\mathbf{K}}_{1,-1}^{(k+1)}(u),\\ \mathbf{G}^{(n+1)}(u)&=\check{\mathbf{K}}_{0,0}^{(k+1)}(u)-\check{\mathbf{K}}_{0,-1}^{(k+1)}(u)\left[\check{\mathbf{K}}_{1,-1}^{(k+1)}(u)\right]^{-1}\check{\mathbf{K}}_{1,0}^{(k+1)}(u).\end{split}$ (S-62)
If $n-M$ is even, i.e. $n-M=2k$ ($k\in\mathbb{Z}_{+}$) the K-matrix $\mathbf{K}^{(M+1)}$ is a $Y(\mathfrak{o}_{4k+1},\mathfrak{o}_{4k+1})$ representation therefore the remaining $\mathbf{G}$ functions are defined in the previous subsection. we have to deform the K-matrix $\mathbf{K}^{(M+1)}$ with (S-58) where we replace $n\to\tilde{n}=2k=n-M$ , $\mathbf{K}\to\mathbf{K}^{(M+1)}$ and $\kappa_{N}\to\kappa_{N-2M}$. The deformed $K$-matrix is $\tilde{\mathbf{K}}^{(1)}$. We also define the nested $K$-matrices $\tilde{\mathbf{K}}^{(s)}$ for $s=1,\dots,k$ with the recursion (S-50) and the $Y^{-}(2)$ $K$-matrices $\mathbf{k}^{(s)}$ the deformed ones $\tilde{\mathbf{k}}^{(s)}$ for $s=1,\dots,k$ with (S-52). The G-operators $\mathbf{G}^{(M+1)},\dots,\mathbf{G}^{(n)}$ are defined with (S-59) for $s=1,\dots,k$. The remaining $G$-operator $\mathbf{G}^{(n+1)}$ is defined as
$\mathbf{G}^{(n+1)}(u)=\mathbf{K}_{0,0}^{(k)}(u)-\sum_{\alpha,\beta=1}^{2}\mathbf{K}_{0,-\alpha}^{(k)}(u)\widehat{\mathbf{K}}_{-\alpha,\beta}^{(k)}(u)\mathbf{K}_{\beta,0}^{(k)}(u).$ (S-63)
@@
In the $\mathfrak{g}=\mathfrak{sp}_{2n}$ case there exists only chiral pair structure. We introduce $n+1$ G-operators $\mathbf{G}^{(k)}(u)$ for $k=1,\dots,n+1$ and we introduce the $F$-operators as (S-23) and the final $\tilde{\mathcal{F}}$ functions defined as
~$\tilde{\mathcal{F}}_{j}^{(k)}(u)=\begin{cases}\mathcal{F}_{j}^{(k)}(-iu+\frac{n-k}{2})\sqrt{\frac{u^{2}}{u^{2}+1/4}},&k=1,\dots,n-1,\\ \mathcal{F}_{j}^{(n)}(-iu)\sqrt{\frac{u^{2}}{u^{2}+1}},&k=n.\end{cases}$ (S-64)
@@
We have a series of $K$-matrices $\mathbf{K}^{(k)}$ for $k=1,\dots,n$ with the recursion (S-40) with the initial condition $\mathbf{K}^{(1)}(u)\equiv\mathbf{K}(u)$. We also define the G-operators $\mathbf{G}^{(k)}$ for $k=1,\dots,n$ with (S-41) and $\mathbf{G}^{(n+1)}$ with (S-42).
@@
We have a series of $K$-matrices $\mathbf{K}^{(s)}$ for $s=1,\dots,m+1$ with the recursion (S-50) and initial condition $\mathbf{K}^{(1)}\equiv\mathbf{K}$. We also define the $Y^{-}(2)$ $K$-matrices $\mathbf{k}^{(s)}$ and the deformed ones $\tilde{\mathbf{k}}^{(s)}$ for $s=1,\dots,m$ with (S-52). We can define the G-operators $\mathbf{G}^{(1)},\dots,\mathbf{G}^{(2m)}$ with (S-37) for $s=1,\dots,m$. For $n>2m$ we also need to define the remaining $G$-operators $\mathbf{G}^{(2m+1)}(u),\mathbf{G}^{(2m+2)}(u),\dots,\mathbf{G}^{(n+1)}(u)$.
The K-matrix $\mathbf{K}^{(2m+1)}$ is a $Y(\mathfrak{sp}_{2n-4m},\mathfrak{sp}_{2n-4m})$ representation. We introduce $k=n-2m$. For $k>0$, we have to deform the K-matrix $\mathbf{K}^{(2m+1)}$ as
$\begin{split}\mathbf{K}_{a,b}^{(2m+1)}(u)&=\sum_{c=-k}^{k}\theta_{a}\theta_{c}\mathbf{L}_{-c,-a}(u-\kappa_{2k})\mathbf{L}_{c,b}(-u),\\ \tilde{\mathbf{K}}_{a,b}^{(1)}(u)&=\mathbf{K}_{a,b}^{(2m+1)}(u)+\sum_{c=1}^{k}\theta_{a}\theta_{c}\mathbf{L}_{-c,-a}(u-\kappa_{2k})\mathbf{L}_{-c,b}(-u),\end{split}$ (S-65)
@@
$\theta_{i}=\begin{cases}+1,&i>0,\\ -1,&i<0.\end{cases}$ (S-66)
@@
$\mathbf{G}^{(2m+s)}(u)=\tilde{\mathbf{K}}_{k-s+1,-k+s-1}^{(s)}(u),$ (S-67)
@@
$\mathbf{G}^{(n+1)}(u)=\tilde{\mathbf{K}}_{-1,1}^{(k)}(u)-\tilde{\mathbf{K}}_{-1,-1}^{(k)}(u)\left[\tilde{\mathbf{K}}_{1,-1}^{(k)}(u)\right]^{-1}\tilde{\mathbf{K}}_{1,1}^{(k)}(u).$ (S-68)
@@
In this section, we demonstrate the application of the overlap formulas on a concrete example.
@@
$H=\sum_{j=1}^{J}\mathbf{P}_{j,j+1}-\frac{1}{2}\mathbf{Q}_{j,j+1},$ (S-69)
@@
$\mathbf{P}=\sum_{i,j=1}^{6}e_{i,j}\otimes e_{j,i},\quad\mathbf{Q}=\sum_{i,j=1}^{6}e_{i,j}\otimes e_{7-i,7-j}.$ (S-70)

### v2

@@ Acknowledgments
$\mathbf{F}^{(n)}(u)=\left[\tilde{\mathbf{K}}_{2,-1}^{(k)}(u)\right]^{-1}\left[\tilde{\mathbf{K}}_{-1,2}^{(k)}(u)-\tilde{\mathbf{K}}_{-1,-1}^{(k)}(u)\left[\tilde{\mathbf{K}}_{2,-1}^{(k)}(u)\right]^{-1}\tilde{\mathbf{K}}_{2,2}^{(k)}(u)\right].$ (S-55)
S-IV.1.2 $Y(\mathfrak{o}_{4k+2},\mathfrak{gl}_{2k+1})$ (achiral)
We define the nested $K$-matrices $\mathbf{K}^{(s)}$ for $s=1,\dots,k$ with the recursion (S-50) and initial condition $\mathbf{K}^{(1)}\equiv\mathbf{K}$. We also define the $Y^{-}(2)$ $K$-matrices $\mathbf{k}^{(s)}$ the deformed ones $\tilde{\mathbf{k}}^{(s)}$ for $s=1,\dots,k$ with (S-53). The G-operators $\mathbf{G}^{(1)},\dots,\mathbf{G}^{(n-1)}$ are defined with (S-37) for $s=1,\dots,k$. The remaining $G$-operator is
$\mathbf{G}^{(n)}(u)=\mathbf{K}_{1,1}^{(k)}(u)-\sum_{\alpha,\beta=2}^{3}\mathbf{K}_{1,-\alpha}^{(k)}(u)\widehat{\mathbf{K}}_{-\alpha,\beta}^{(k)}(u)\mathbf{K}_{\beta,1}^{(k)}(u).$ (S-56)
@@
We have a series of $K$-matrices $\mathbf{K}^{(s)}$ for $s=1,\dots,n-1$ with the recursion (S-40). We also define the G-operators $\mathbf{G}^{(s)}$ for $s=1,\dots,n-1$ with (S-41). We also need to define the remaining $G$-operators $\mathbf{G}^{(n)}(u),\mathbf{G}^{(n+1)}(u)$ as
$\begin{split}\mathbf{G}^{(n)}(u)&=\mathbf{K}_{1,-1}^{(n-1)}(u)-\mathbf{K}_{1,-2}^{(n-1)}(u)\left[\mathbf{K}_{2,-2}^{(n-1)}(u)\right]^{-1}\mathbf{K}_{2,-1}^{(n-1)}(u),\\ \mathbf{G}^{(n+1)}(u)&=\mathbf{K}_{-1,1}^{(n-1)}(u)-\mathbf{K}_{-1,-2}^{(n-1)}(u)\left[\mathbf{K}_{2,-2}^{(n-1)}(u)\right]^{-1}\mathbf{K}_{2,1}^{(n-1)}(u).\end{split}$ (S-57)
@@
$\mathbf{G}^{(n)}(u)=\mathbf{K}_{1,1}^{(n-1)}(u)-\mathbf{K}_{1,-2}^{(n-1)}(u)\left[\mathbf{K}_{2,-2}^{(n-1)}(u)\right]^{-1}\mathbf{K}_{2,1}^{(n-1)}(u).$ (S-58)
@@
We continue with general $n$ and $M$. We have a series of $K$-matrices $\mathbf{K}^{(s)}$ for $s=1,\dots,M+1$ with the recursion (S-40). We also define the G-operators $\mathbf{G}^{(s)}$ for $s=1,\dots,M$ with (S-41). We also need to define the remaining $G$-operators $\mathbf{G}^{(M+1)}(u),\mathbf{G}^{(M+2)}(u),\dots,\mathbf{G}^{(n+1)}(u)$ for the chiral and $\mathbf{G}^{(M+1)}(u),\mathbf{G}^{(M+2)}(u),\dots,\mathbf{G}^{(n)}(u)$ for achiral cases. When $n-M$ is even the pair structure is chiral and the pair structure is odd when $n-M$ is odd.
@@
$\begin{split}\mathbf{K}_{a,b}^{(M+1)}(u)&=\sum_{c=-\tilde{n}}^{\tilde{n}}\mathbf{L}_{-c,-a}(u-\kappa_{N-2M})\mathbf{L}_{c,b}(-u),\\ \tilde{\mathbf{K}}_{a,b}^{(1)}(u)&=\mathbf{K}_{a,b}^{(M+1)}(u)+\\ &+u\sum_{l=1}^{k}\left(\mathbf{L}_{-(\tilde{n}-2l+1),-a}(u-\kappa_{N-2M})\mathbf{L}_{-\tilde{n}+2l-2,b}(-u)-\mathbf{L}_{-(\tilde{n}-2l+2),-a}(u-\kappa_{N-2M})\mathbf{L}_{-\tilde{n}+2l-1,b}(-u)\right),\end{split}$ (S-59)
where $\tilde{n}=2k=n-M$. We continue with the series of the nested $K$-matrices $\tilde{\mathbf{K}}_{a,b}^{(s)}(u)$ for $s=2,\dots,k$ as (S-50) and define the $Y^{-}(2)$ $K$-matrices $\mathbf{k}^{(s)}$ and the deformed ones $\tilde{\mathbf{k}}^{(s)}$ for $s=1,\dots,k$ with (S-53) (we replace $n\to\tilde{n}$ , $\mathbf{K}^{(s)}\to\tilde{\mathbf{K}}^{(s)}$ and $\kappa_{N-4s+4}\to\kappa_{N-2M-4s+4}$). The G-operators $\mathbf{G}^{(M+1)},\dots,\mathbf{G}^{(n)}$ are
$\begin{split}\mathbf{G}^{(M+2s-1)}(u)&=\tilde{\mathbf{k}}_{1,1}^{(s)}(u),\\ \mathbf{G}^{(M+2s)}(u)&=\tilde{\mathbf{k}}_{2,2}^{(s)}(u)-\tilde{\mathbf{k}}_{2,1}^{(s)}(u)\left[\tilde{\mathbf{k}}_{1,1}^{(s)}(u)\right]^{-1}\tilde{\mathbf{k}}_{1,2}^{(s)}(u),\end{split}$ (S-60)
for $s=1,\dots,k$. The F-operators are defined as (S-54) and (S-55).
If $n-M$ is odd, i.e. $n-M=2k+1$ ($k\in\mathbb{Z}_{+}$) we have to deform the K-matrix $\mathbf{K}^{(M+1)}$ as (S-59) where $\tilde{n}=2k+1=n-M$. We continue with the series of the nested $K$-matrices $\tilde{\mathbf{K}}_{a,b}^{(s)}(u)$ for $s=2,\dots,k$ as (S-50). We define the $Y^{-}(2)$ $K$-matrices $\mathbf{k}^{(s)}$ and the deformed ones $\tilde{\mathbf{k}}^{(s)}$ for $s=1,\dots,k$ with (S-53) (we replace $n\to\tilde{n}$ , $\mathbf{K}^{(s)}\to\tilde{\mathbf{K}}^{(s)}$ and $\kappa_{N-4s+4}\to\kappa_{N-2M-4s+4}$). The G-operators $\mathbf{G}^{(M+1)},\dots,\mathbf{G}^{(n-1)}$ are defined with (S-60) for $s=1,\dots,k$. The remaining $G$-operator $\mathbf{G}^{(n)}$ is defined as (S-56).
@@
In the $\mathfrak{g}=\mathfrak{o}_{2n+1}$ case there exists only chiral pair structure. We introduce $n+1$ $G$-operators $\mathbf{G}^{(k)}(u)$ for $k=1,\dots,n+1$ and we introduce the $F$-operators as (S-22). The final $\tilde{\mathcal{F}}$ functions are defined as
~$\tilde{\mathcal{F}}_{j}^{(k)}(u)=\begin{cases}\mathcal{F}_{j}^{(k)}(iu-\frac{1}{4}+\frac{n-k}{2})\sqrt{\frac{u^{2}}{u^{2}+1/4}},&k=1,\dots,n-1,\\ \mathcal{F}_{j}^{(n)}(iu-\frac{1}{4})\frac{u+i/4}{u+i/2}\sqrt{\frac{u^{2}}{u^{2}+1/16}},&k=n.\end{cases}$ (S-61)
@@
For $Y(\mathfrak{o}_{2n+1},\mathfrak{o}_{n}\oplus\mathfrak{o}_{n+1})$ we have a series of $K$-matrices $\mathbf{K}^{(k)}$ for $k=1,\dots,n$ with the recursion (S-40) and initial condition $\mathbf{K}^{(1)}(u)\equiv\mathbf{K}(u)$. We also define the G-operators $\mathbf{G}^{(k)}$ for $k=1,\dots,n$ with (S-41) and $\mathbf{G}^{(n+1)}$ with (S-42).
@@
For general $M$ we have a series of $K$-matrices $\mathbf{K}^{(s)}$ for $s=1,\dots,M+1$ with the recursion (S-40) and initial condition $\mathbf{K}^{(1)}(u)\equiv\mathbf{K}(u)$. We also define the G-operators $\mathbf{G}^{(s)}$ for $s=1,\dots,M$ with (S-41). We also need to define the remaining $G$-operators $\mathbf{G}^{(M+1)}(u),\mathbf{G}^{(M+2)}(u),\dots,\mathbf{G}^{(n+1)}(u)$. Their definitions depend on whether $n-M$ is even or odd.
If $n-M$ is odd, i.e. $n-M=2k+1$ ($k\in\mathbb{N}$) we have to deform the K-matrix $\mathbf{K}^{(M+1)}$ with (S-59) where $\tilde{n}=2k+1=n-M$. We define the nested $K$-matrices $\tilde{\mathbf{K}}_{a,b}^{(s)}(u)$ for $s=2,\dots,k+1$ with (S-50), the $Y^{-}(2)$ $K$-matrices $\mathbf{k}^{(s)}$ and the deformed ones $\tilde{\mathbf{k}}^{(s)}$ for $s=1,\dots,k$ with (S-53) (we replace $n\to\tilde{n}$ , $\mathbf{K}^{(s)}\to\tilde{\mathbf{K}}^{(s)}$ and $\kappa_{N-4s+4}\to\kappa_{N-2M-4s+4}$). We can also define the G-operators $\mathbf{G}^{(r)}$ for $r=M+1,\dots,n-1$ as (S-60) for $s=1,\dots,k$. We also have to deform the K-matrix $\tilde{\mathbf{K}}_{a,b}^{(k+1)}(u)$ as
$\begin{split}\tilde{\mathbf{K}}_{a,b}^{(k+1)}(u)&=\sum_{c=-1}^{1}\mathbf{L}_{-c,-a}(u-\kappa_{3})\mathbf{L}_{c,b}(-u),\\ \check{\mathbf{K}}_{a,b}^{(k+1)}(u)&=\tilde{\mathbf{K}}_{a,b}^{(k+1)}(u)-\frac{1}{2}u(u+\frac{1}{4})\mathbf{L}_{-1,-a}(u-\kappa_{3})\mathbf{L}_{-1,b}(-u)\\ &+u\left(\mathbf{L}_{0,-a}(u-\kappa_{3})\mathbf{L}_{-1,b}(-u)-\mathbf{L}_{-1,-a}(u-\kappa_{3})\mathbf{L}_{0,b}(-u)\right).\end{split}$ (S-62)
and the remaining $G$-operators $\mathbf{G}^{(n)}$, $\mathbf{G}^{(n+1)}$ are
$\begin{split}\mathbf{G}^{(n)}(u)&=\check{\mathbf{K}}_{1,-1}^{(k+1)}(u),\\ \mathbf{G}^{(n+1)}(u)&=\check{\mathbf{K}}_{0,0}^{(k+1)}(u)-\check{\mathbf{K}}_{0,-1}^{(k+1)}(u)\left[\check{\mathbf{K}}_{1,-1}^{(k+1)}(u)\right]^{-1}\check{\mathbf{K}}_{1,0}^{(k+1)}(u).\end{split}$ (S-63)
If $n-M$ is even, i.e. $n-M=2k$ ($k\in\mathbb{Z}_{+}$), we have to deform the K-matrix $\mathbf{K}^{(M+1)}$ with (S-59) where we replace $n\to\tilde{n}=2k=n-M$ , $\mathbf{K}\to\mathbf{K}^{(M+1)}$ and $\kappa_{N}\to\kappa_{N-2M}$. We also define the nested $K$-matrices $\tilde{\mathbf{K}}^{(s)}$ for $s=1,\dots,k$ with the recursion (S-50) and the $Y^{-}(2)$ $K$-matrices $\mathbf{k}^{(s)}$ the deformed ones $\tilde{\mathbf{k}}^{(s)}$ for $s=1,\dots,k$ with (S-53). The G-operators $\mathbf{G}^{(M+1)},\dots,\mathbf{G}^{(n)}$ are defined with (S-60) for $s=1,\dots,k$. The remaining $G$-operator $\mathbf{G}^{(n+1)}$ is defined as
$\mathbf{G}^{(n+1)}(u)=\mathbf{K}_{0,0}^{(k)}(u)-\sum_{\alpha,\beta=1}^{2}\mathbf{K}_{0,-\alpha}^{(k)}(u)\widehat{\mathbf{K}}_{-\alpha,\beta}^{(k)}(u)\mathbf{K}_{\beta,0}^{(k)}(u).$ (S-64)
@@
In the $\mathfrak{g}=\mathfrak{sp}_{2n}$ case there exists only chiral pair structure. We introduce $n+1$ G-operators $\mathbf{G}^{(k)}(u)$ for $k=1,\dots,n+1$ and we introduce the $F$-operators as (S-22) and the final $\tilde{\mathcal{F}}$ functions are defined as
~$\tilde{\mathcal{F}}_{j}^{(k)}(u)=\begin{cases}\mathcal{F}_{j}^{(k)}(iu+\frac{n-k}{2})\sqrt{\frac{u^{2}}{u^{2}+1/4}},&k=1,\dots,n-1,\\ \mathcal{F}_{j}^{(n)}(iu)\sqrt{\frac{u^{2}}{u^{2}+1}},&k=n.\end{cases}$ (S-65)
@@
We have a series of $K$-matrices $\mathbf{K}^{(k)}$ for $k=1,\dots,n$ with the recursion (S-40) and initial condition $\mathbf{K}^{(1)}(u)\equiv\mathbf{K}(u)$. We also define the G-operators $\mathbf{G}^{(k)}$ for $k=1,\dots,n$ with (S-41) and $\mathbf{G}^{(n+1)}$ with (S-42).
@@
We have a series of $K$-matrices $\mathbf{K}^{(s)}$ for $s=1,\dots,m+1$ with the recursion (S-50) and initial condition $\mathbf{K}^{(1)}\equiv\mathbf{K}$. We also define the $Y^{-}(2)$ $K$-matrices $\mathbf{k}^{(s)}$ and the deformed ones $\tilde{\mathbf{k}}^{(s)}$ for $s=1,\dots,m$ with (S-53). We can define the G-operators $\mathbf{G}^{(1)},\dots,\mathbf{G}^{(2m)}$ with (S-37) for $s=1,\dots,m$. For $n>2m$ we also need to define the remaining $G$-operators $\mathbf{G}^{(2m+1)}(u),\mathbf{G}^{(2m+2)}(u),\dots,\mathbf{G}^{(n+1)}(u)$.
For $k=n-2m>0$, we have to deform the K-matrix $\mathbf{K}^{(2m+1)}$ as
$\begin{split}\mathbf{K}_{a,b}^{(2m+1)}(u)&=\sum_{c=-k}^{k}\theta_{a}\theta_{c}\mathbf{L}_{-c,-a}(u-\kappa_{2k})\mathbf{L}_{c,b}(-u),\\ \tilde{\mathbf{K}}_{a,b}^{(1)}(u)&=\mathbf{K}_{a,b}^{(2m+1)}(u)+\sum_{c=1}^{k}\theta_{a}\theta_{c}\mathbf{L}_{-c,-a}(u-\kappa_{2k})\mathbf{L}_{-c,b}(-u),\end{split}$ (S-66)
@@
$\theta_{i}=\begin{cases}+1,&i>0,\\ -1,&i<0.\end{cases}$ (S-67)
@@
$\mathbf{G}^{(2m+s)}(u)=\tilde{\mathbf{K}}_{k-s+1,-k+s-1}^{(s)}(u),$ (S-68)
@@
$\mathbf{G}^{(n+1)}(u)=\tilde{\mathbf{K}}_{-1,1}^{(k)}(u)-\tilde{\mathbf{K}}_{-1,-1}^{(k)}(u)\left[\tilde{\mathbf{K}}_{1,-1}^{(k)}(u)\right]^{-1}\tilde{\mathbf{K}}_{1,1}^{(k)}(u).$ (S-69)
@@
In this section, we demonstrate the application of the overlap formulas on concrete examples.
S-VII.1 $SO(6)$ spin chain
@@
$H=\sum_{j=1}^{J}\mathbf{P}_{j,j+1}-\frac{1}{2}\mathbf{Q}_{j,j+1},$ (S-70)
@@
$\mathbf{P}=\sum_{i,j=1}^{6}e_{i,j}\otimes e_{j,i},\quad\mathbf{Q}=\sum_{i,j=1}^{6}e_{i,j}\otimes e_{7-i,7-j}.$ (S-71)